OpenStudy (anonymous):

Find the probability. An IRS auditor randomly selects 3 tax returns (without replacement) from 47 returns of which 5 contain errors. What is the probability that she selects none of those containing errors? A. 0.0012 B. 0.0006 C. 0.708 D. 0.7136

OpenStudy (anonymous):

of the 47, 5 contain errors and so 42 do not. probability of selecting 3 that do not is \[\frac{42}{47}\times \frac{41}{46}\times \frac{40}{45}\] hope it is clear where the numbers come from

OpenStudy (anonymous):

first one doesn't AND second one doesn't given first one doesn't AND third one doesn't given the first two don't

OpenStudy (anonymous):

so what would my anwser be?